Show that only odd integral multiples of
fundamental frequency are excited in a
closed pipe.
1
Expert's answer
2019-05-03T11:23:51-0400

For a closed pipe the node is at closed end and anti-nodes at open end.

Let, length of the closed pipe is l.

For the fundamental mode ,corresponding wavelength is given by,"\lambda_1 = 4l"

Now, if "n_1" ​ is fundamental frequency and v is velocity of sound in air then, "n_1=\dfrac{v}{\lambda_1} = \dfrac{v}{4l}"

Again for first overtone ,"\lambda_3 = \dfrac{4l}{3}"

So, frequency "n_3 = \dfrac{v}{\lambda_3} = \dfrac{3v}{4l} = 3n_1" [As, "n_1 = \dfrac{v}{4l}" ]

 

Similarly, "n_5 = 5n_1"

 

So, frequency of the m-th overtone is (2m +1) "n_1" ​ . [where, "n_1" ​ is the fundamental frequency]

 

So it is shows that in a closed pipe only odd harmonics are produced.

 

Answer: In a closed pipe odd harmonics are produced.